# Injecting Domain-Specific Knowledge into Large Language Models: A Comprehensive Survey

### v1

@@ 4.3 Materials
In contrast to biomedicine, where significant efforts have been devoted to static knowledge embedding due to the availability of extensive corpora, research in materials and chemistry has primarily focused on utilizing task-related tools that align with the dynamic knowledge injection paradigm.
For instance,  Xie et al. (2024) demonstrated how Darwin 1.5 leverages natural language inputs and a two-stage training strategy to achieve significant improvements in materials discovery and design tasks. Bran et al. (2023) introduced ChemCrow, a framework that augments LLMs with chemistry-expert-designed tools for downstream tasks such as organic synthesis and drug discovery. There are also studies on prompt optimization Tang and others (2025), demonstrating that better-designed planning prompts can effectively utilize the model’s internal knowledge to orchestrate complex tasks. This approach capitalizes on the planning and execution capabilities of multiple LLMs to achieve autonomy in chemical experimentation.
More recently, there has been growing interest in exploring static knowledge embedding and modular knowledge adapters within the chemistry domain. For example, Chen and others (2024) curated a QA dataset to fine-tune pretrained models like BERT and LLMs such as Llama, aiming to enhance their performance in chemistry-related tasks. Similarly, Xie et al. (2024) introduced Darwin 1.5, an open-source large language model tailored for materials science.
@@ 4.4 Human-Centered Science
The last domain we introduce is human-centered science, which encompasses a wide range of applications such as psychological counseling, financial forecasting, social behavior prediction, and legal reasoning. All these domains center on understanding and interacting with human needs, behaviors, and decision-making processes.
In mental health, datasets like PsyQA (Sun et al., 2021) provide a foundation for training models in psychological counseling scenarios. For instance, SoulChat (Chen et al., 2023), a model fine-tuned on over 100,000 long-text counseling sessions using static knowledge embedding, is designed for empathic conversations. Similarly, MeChat (Qiu et al., 2023) employs dynamic knowledge injection to adapt to real-time inputs, significantly enhancing its emotional support capabilities. These advancements demonstrate the potential of human-centered science in addressing complex, real-world challenges through personalized and context-aware solutions.
In the education domain, LLMs have shown immense potential in addressing challenges such as personalized learning, curriculum alignment, and interactive teaching. Personalized learning, for example, requires models to adapt to individual needs, providing tailored feedback and emotional support. EduChat (Dan et al., 2023) tackles this by leveraging educational theories from psychology and pedagogy through static knowledge embedding, enabling tasks like open Q&A, composition correction, and emotional support. Similarly, QiaoBan (Weixiang et al., 2023) focuses on child-centered education by using prompt optimization to adapt the model’s behavior based on child psychology and emotional well-being, catering specifically to young learners. Domain-specific education and interactive teaching have also seen advancements through LLMs. CyberQ (Agrawal et al., 2024) blends static knowledge embedding and dynamic knowledge injection via AISecKG (Agrawal, 2023), generating Q&A based on cybersecurity best practices. Interactive teaching, on the other hand, benefits from models like SocraticLM (Liu et al., 2024c), which employs adapters fine-tuned on the SocraTeach dataset to engage students in critical thinking and problem-solving.
For social sciences, models like SocialLLM (Jiang and Ferrara, 2023) combine static knowledge embedding and dynamic knowledge injection to analyze human behavior in social networks. Adapters facilitate large-scale data integration while prompt optimization guides the model to focus on specific social behavior patterns. Models like FPS (Liu et al., 2024e) and FUSE (Liu et al., 2024f) use prompt optimization to simulate the spread and evolution of fake news in social networks, helping understand misinformation’s impact.
A summary of the mainstream models and their information is provided in Table 3. More models across various domains can be found at: Survey-official-repo.
@@ 5.1 Knowledge Injection Framework
In this section, we provide a detailed introduction to four open-source frameworks categorized under different knowledge injection methods to facilitate understanding and application: KnowGPT (Zhang et al., 2024c) for Dynamic Knowledge Injection, StructTuning (Liu et al., 2024d) for Static Knowledge Embedding, K-Adapter (Wang et al., 2021) for Modular Knowledge Adapters, and SelfLift (Cheng et al., 2024) for Prompt Optimization.
@@
Finally, SelfLift iteratively employs a retrieval-augmented generator to create an unbounded memory pool and uses a memory selector to choose one output as memory for the subsequent generation round. This is an excellent demonstration of prompt optimization, where the model’s outputs are dynamically refined and reused to enhance its overall performance and coherence in subsequent tasks.
5.2 Datasets and Benchmarks
We summarize commonly used datasets or benchmarks for domain-specific LLM study in Table 3, observing significant variation in dataset richness across domains. Biomedicine boasts numerous high-quality datasets, such as PubMed, PubMedQA (Jin et al., 2019), and BioASQ Tsatsaronis and others (2012), that support tasks such as question answering and clinical summarization. In contrast, materials and chemistry have more limited resources, and datasets like USPTO and Enzymes focus on chemical reactions. Miscellaneous datasets are scattered across other domains, such as PsyQA and SmileChat in mental health, SocraTeach, and Children’s emotional education dialogue data dataset in education. This diversity underscores the effort to tailor LLMs to specialized fields while emphasizing the need for broader curation of benchmarks in underrepresented domains.
5.3 Performance Comparison of Domain-specific LLM and General-domain LLM
Model Domain Type Size MedQA PubMedQA MedMCQA
Med-Gemini Specific 91.1 - -
@@ 5.3 Performance Comparison of Domain-specific LLM and General-domain LLM
Med-PaLM 2 Specific 85.4 81.8 72.3
PMC-LLaMA Specific 13B 56.3 77.9 56.0
BioMedLM Specific 2.7B 50.3 74.4
Llama 2 General 70B 43.7 74.3 35.0
Galactica General 120B 44.4 77.6 52.9
Table 4: Performance comparison of domain-specific and general-domain model performance on medical benchmarks.
Since there are also general-domain strong LLMs, it is necessary to discuss the comparison between domain-specific LLMs and general ones to determine if a specific knowledge injection process is essential. Here, we take the biomedical domain as an example due to significant research efforts in this field, as shown in Table 4. The results are collected from corresponding papers or paperswithcode.com.
First, we can observe that closed-source LLMs are currently the most effective models, while the performance gap between general-domain and domain-specific LLMs is relatively narrow. For example, both GPT-4 and Med-Gemini (Saab et al., 2024) achieve outstanding performance, with scores higher than 90 on the MedQA dataset. However, due to the lack of transparency in closed-source LLMs, open-source LLM efforts should not be overlooked. In this field, domain-specific LLMs often outperform general-domain models. For instance, PMC LLaMA-13B outperforms LLaMA2-70B by more than 10 points on the MedQA dataset. This demonstrates the value of domain-specific LLMs in achieving superior performance on specialized tasks. While general-domain models can deliver strong results, incorporating domain-specific knowledge allows for significant improvements, particularly in open-source initiatives. This underscores the importance of investing in domain-specific LLMs to address the unique challenges of specialized fields.
@@ 6 Challenges and Opportunities
6.1 Integrated Knowledge Consistency
Knowledge injection allows LLMs to incorporate and integrate different domain-specific knowledge. However, retrieved knowledge may conflict with the model’s pre-trained representations or other retrieved facts, leading to inconsistencies in outputs (Xu et al., 2024a). For example, in healthcare or legal analysis, conflicting treatment protocols or contradictory legal precedents could arise, resulting in unreliable decisions and undermining the system’s trustworthiness.
To address this, future research must focus on detecting inconsistencies, resolving conflicts, and maintaining consistency in integrated knowledge. Conflicts can be addressed by prioritizing reliable sources, applying domain-specific rules, or using ensemble techniques to balance multiple perspectives. Alignment algorithms and validation modules can further ensure that retrieved knowledge aligns with the model’s reasoning processes and is reliable before influencing outputs. These efforts are essential to enhance the reliability and applicability of knowledge-enhanced LLMs in complex, high-stakes domains.
6.2 Cross-Domain Knowledge Transfer
Cross-domain knowledge transfer involves equipping LLMs with the ability to generalize knowledge across diverse and distinct fields. While this significantly expands their applicability, it also introduces challenges due to the complexity and diversity of domain-specific terminologies, ontologies, and reasoning patterns. For example, transferring knowledge from chemistry to healthcare might require reconciling differing data structures and reasoning frameworks.
Overcoming these challenges necessitates advancements in modular knowledge representation and transfer learning techniques. Future efforts could explore hybrid approaches that blend static embeddings with dynamic retrieval, enabling LLMs to adapt knowledge flexibly across domains without compromising depth. Additionally, creating standardized cross-domain benchmarks and datasets could facilitate systematic evaluation and foster innovation in multi-domain knowledge transfer methodologies.
@@ 7 Conclusion
LLMs enhanced by domain-specific knowledge have shown remarkable potential and garnered increasing research interest. This survey systematically reviews LLM knowledge injection systems, exploring knowledge representation methods, integration strategies, and mechanisms for preserving model generality. We also summarize applications across biomedicine, chemistry, and computational social science domains. By highlighting standard datasets, benchmarks, challenges, and future opportunities, we aim to provide a valuable resource that inspires further exploration of knowledge-enhanced LLMs for domain-specific challenges.

### v2

@@ 4.3 Materials
In contrast to the biomedical domain, the field of materials science and chemistry has largely focused on static knowledge embedding. Many models rely on domain-specific corpora to fine-tune general models for improved task performance. Darwin 1.5 (Xie et al., 2024) adopts a two-stage training strategy using natural language inputs to enhance performance in materials discovery. ScholarChemQA (Chen et al., 2024) constructs a chemistry QA dataset to fine-tune BERT and LLaMA, improving chemical reasoning. Recently, some efforts have begun to explore dynamic knowledge integration. ChemCrow (Bran et al., 2023) augments LLMs with chemistry tools for applications like synthesis and drug discovery. ChemAgent (Tang et al., 2025b) shows that well-designed planning prompts can guide models through complex execution tasks by leveraging internal reasoning.
While still in early stages, the field is transitioning from static embedding toward interactive and tool-augmented reasoning, indicating strong potential for future developments
@@ 4.4 Human-Centered Science
Human-centered science focuses on understanding and assisting human behaviors, needs, and decisions. This interdisciplinary domain includes mental health, education, social behavior prediction, and legal reasoning—each benefiting from personalized and context-aware LLMs.
In mental health, datasets like PsyQA (Sun et al., 2021) provide a foundation for training models in psychological counseling scenarios. SoulChat (Chen et al., 2023), a model fine-tuned on over 100,000 long-text counseling sessions using static knowledge embedding, is designed for empathic conversations. In contrast, MeChat (Qiu et al., 2023) employs dynamic knowledge injection to adapt to real-time inputs, enhancing its emotional support capabilities. These advancements demonstrate the potential of human-centered science in addressing complex, real-world challenges through personalized and context-aware solutions.
In the education domain, LLMs have shown immense potential in addressing challenges such as personalized learning, curriculum alignment, and interactive teaching. Personalized learning, for example, requires models to adapt to individual needs, providing tailored feedback and emotional support. EduChat (Dan et al., 2023) applies psychological and pedagogical theories via static knowledge embedding to support tasks like Q&A, writing feedback, and emotional guidance. Similarly, QiaoBan (Weixiang et al., 2023) uses prompt optimization to tailor model behavior to children’s psychological and emotional needs. Domain-specific education and interactive teaching have also seen advancements. CyberQ (Agrawal et al., 2024) blends static knowledge embedding and dynamic knowledge injection via AISecKG (Agrawal, 2023), generating Q&A based on cybersecurity best practices. Interactive teaching, on the other hand, benefits from models like SocraticLM (Liu et al., 2024c), which employs adapters fine-tuned on the SocraTeach dataset to engage students in critical thinking and problem-solving.
For social sciences, models like SocialLLM (Jiang and Ferrara, 2023) combine static knowledge embedding and dynamic knowledge injection to analyze human behavior in social networks. Models like SSF (Wang et al., 2024a) FPS (Liu et al., 2024e) and FUSE (Liu et al., 2025b) use prompt optimization to simulate the spread and evolution of fake news in social networks, helping understand misinformation’s impact. In addition, Liu et al. (2025a) adopts a multi-agent approach to synthesize data, which is then used as knowledge to enhance language models for detecting fake news. A summary of the mainstream models and their information is provided in Table 3. More models across various domains can be found at: official-repo.
@@ 5.1 Knowledge Injection Framework
In this section, we provide a detailed introduction to four open-source frameworks categorized under different knowledge injection methods to facilitate understanding and application: KnowGPT (Zhang et al., 2024d) for Dynamic Knowledge Injection, StructTuning (Liu et al., 2024d) for Static Knowledge Embedding, K-Adapter (Wang et al., 2021) for Modular Knowledge Adapters, and SelfLift (Cheng et al., 2024) for Prompt Optimization.
@@
Finally, SelfLift iteratively employs a retrieval-augmented generator to create an unbounded memory pool and uses a memory selector to choose one output as memory for the subsequent generation round. This is a good demonstration of prompt optimization, where the model’s outputs are dynamically refined and reused to enhance its overall performance and coherence in subsequent tasks.
5.2 Knowledge Source
We summarize commonly used knowledge sources for domain-specific LLMs in Table 3, referring to datasets that provide the external knowledge used in various injection methods—including training corpora for static embedding or adapter tuning, and retrieval or prompt design resources for dynamic knowledge injection. Biomedicine includes numerous high-quality datasets, such as PubMed, PubMedQA (Jin et al., 2019), and BioASQ Tsatsaronis et al. (2012), which support tasks such as question answering and clinical summarization. In contrast, materials and chemistry have more limited resources, and datasets like USPTO and Enzymes focus on chemical reactions. Miscellaneous datasets are scattered across other domains, such as PsyQA and SmileChat in mental health, SocraTeach, and Children’s emotional education dialogue data dataset in education. This diversity underscores the effort to tailor LLMs to specialized fields while emphasizing the need for broader curation of benchmarks in underrepresented domains.
5.3 Performance Comparison of 4 Paradigms
Model Category MedQA PubMedQA MedMCQA
GPT-4 (Medprompt) Prompt Optimization 90.2 82.0 79.1
@@ 5.3 Performance Comparison of 4 Paradigms
Med-PaLM 2 Static Knowledge 85.4 81.8 72.3
Flan-PaLM (3-shot) Dynamic Knowledge 67.6 79.0 57.6
PMC-LLaMA Static Knowledge 56.3 77.9 56.0
BioMedLM Static Knowledge 50.3 74.4
LLaMA (MedAdapter) Knowledge Adapters 37.4 63.6 32.0
Table 4: Model performance across four knowledge paradigms on medical benchmarks.
To compare knowledge injection paradigms in a practical setting, we focus on the biomedical domain due to its popularity and the availability of benchmarks such as MedQA, PubMedQA, and MedMCQA, as shown in Table 4. Although the models differ in architecture, we align backbones when possible. For example, both PMC-LLaMA and MedAdapter use LLaMA-13B. SOTA models like GPT-4 are closed-source, making prompt optimization the only feasible adaptation strategy. Despite no domain-specific training, GPT-4 with Medprompt achieves strong performance, showing the effectiveness of prompt methods for closed models. Among open models, MedAdapter underperforms compared to PMC-LLaMA, suggesting that full fine-tuning may outperform adapter-based methods for some tasks. Performance differences across paradigms also highlight the importance of pretraining corpus and task alignment, particularly in static injection approaches. Furthermore, in Appendix A, we also compare knowledge injection paradigms in the finance domain and obtain similar conclusions to those in the medical domain.
@@ 6 Challenges and Opportunities
Integrated Knowledge Consistency. Knowledge injection allows LLMs to incorporate and integrate different domain-specific knowledge. However, retrieved knowledge may conflict with the model’s pre-trained representations or other retrieved facts, leading to inconsistencies in outputs (Xu et al., 2024b). For example, in healthcare or legal analysis, conflicting treatment protocols or contradictory legal precedents could arise Dayton (2012), resulting in unreliable decisions and undermining the system’s trustworthiness. To address this, future research must focus on detecting inconsistencies, resolving conflicts, and maintaining consistency in integrated knowledge. Conflicts can be addressed by prioritizing reliable sources, applying domain-specific rules, or using ensemble techniques to balance multiple perspectives. Alignment and validation modules help ensure retrieved knowledge fits the model’s reasoning.
Cross-Domain Knowledge Transfer. Cross-domain knowledge transfer involves equipping LLMs with the ability to generalize knowledge across diverse and distinct fields Li et al. (2025b). While this significantly expands their applicability, it also introduces challenges due to the complexity and diversity of domain-specific terminologies, ontologies, and reasoning patterns Montero et al. (2004). For example, transferring knowledge from chemistry to healthcare might require reconciling differing data structures and reasoning frameworks (Schroeder et al., 2018). Overcoming these challenges requires advancements in modular knowledge representation and transfer learning techniques. Future efforts could explore hybrid approaches that blend static embeddings with dynamic retrieval, enabling LLMs to adapt knowledge flexibly across domains without compromising depth. Additionally, standardized cross-domain benchmarks can enable consistent evaluation and drive innovation in knowledge transfer methods. We provide more discussions in Appendix B.
@@ 7 Conclusion
LLMs enhanced by domain-specific knowledge have shown remarkable potential and garnered increasing research interest. This survey systematically reviews LLM knowledge injection systems, exploring knowledge representation methods, integration strategies, and mechanisms for preserving model generality. We also summarize applications across biomedicine, chemistry, and computational social science domains. By highlighting standard datasets, benchmarks, challenges, and future opportunities, we aim to provide a valuable resource that inspires exploration of knowledge-enhanced LLMs for domain-specific challenges.
Limitation
Despite providing a comprehensive review of current methods and applications for domain-specific knowledge injection in LLMs, this survey has certain limitations. First, while we strive to cover several key domains such as finance, biomedicine, and materials science, some less-studied or emerging areas (for example, low-resource languages, cross-cultural education, and niche disciplines) receive relatively limited attention. Second, our focus is primarily on summarizing methodological principles and representative models from existing literature. Due to substantial variation in model architectures, application domains, training data, and evaluation protocols, we were only able to conduct targeted comparisons under controlled conditions within the biomedical domain, using commonly adopted datasets. A more systematic and broad-based empirical evaluation across methods remains an important direction for future work. Nevertheless, we hope this survey serves as a useful reference and provides a clear roadmap for ongoing research in knowledge-enhanced LLMs.
Acknowledgement
We would like to thank the anonymous reviewers for their constructive comments. The work was supported by Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) through grant award 8481000078.
Appendix A Performance Comparison of 4 Paradigms
Model Category FPB FiQA-SA TFNS
GPT-4 General 83.3 63.0 80.8
GPT-4(finetune) Static Knowledge 87.8 88.7 88.3
FinGPT Knowledge Adapters 88.2 87.4 90.3
FinBERT Static Knowledge 88.0 59.6 73.3
BloombergGPT Static Knowledge 51.1 75.1 -
Llama2-7B General 39.0 80.0 29.6
FLANG Static Knowledge 91.9 3.4 -
Table 5: Performance comparison of representative models under the four knowledge injection paradigms on financial benchmarks.
To systematically compare the effectiveness of different knowledge injection paradigms in practical settings, we focus on two representative domains: biomedicine and finance.
In the biomedical domain, which is widely studied and rich in benchmark datasets (e.g., MedQA, PubMedQA, and MedMCQA), we evaluate the performance of various models (see Table 4). Although the models differ in architecture, we align their backbones where possible—for example, both PMC-LLaMA and MedAdapter use LLaMA-13B. For closed-source models like GPT-4, prompt engineering is the only feasible adaptation strategy. Nevertheless, GPT-4 combined with Medprompt achieves strong performance, demonstrating the effectiveness of prompt-based knowledge injection. Among open models, MedAdapter underperforms compared to fully fine-tuned models such as PMC-LLaMA, suggesting that full fine-tuning may be more effective than adapter-based methods for certain tasks. Models with static knowledge (e.g., MedBERT) show substantial variance across tasks, underscoring the importance of alignment between pretraining corpora and downstream objectives.
In Table 5, we extend this comparison to the financial domain, evaluating models on benchmarks such as FPBMalo et al. (2014), FiQA-SAMaia et al. (2018), and TFNSEl-Haj et al. (2020). The findings closely mirror those in the biomedical domain. Finetuned GPT-4 consistently outperforms others, confirming the value of injecting domain-specific knowledge into general-purpose LLMs. Static knowledge models like FinBERT and FLANG perform well on certain tasks but show significant variability, again emphasizing the crucial role of corpus-task alignment. FinGPT, which adopts lightweight adapter-based knowledge injection, achieves competitive performance while maintaining adaptability. In contrast, LLaMA2-7B lags behind across most tasks, reinforcing the necessity of targeted knowledge injection for domain-intensive applications. The consistency of observations across both domains suggests that the effectiveness of knowledge injection depends on a careful balance of architectural design, adaptation strategy, and corpus alignment to support complex, high-stakes tasks.
Appendix B Detailed Discussions on Challenges & Opportunities
B.1 Integrated Knowledge Consistency
While knowledge injection empowers LLMs to reason with external facts, it introduces a crucial consistency problem: injected knowledge may contradict either the model’s internal representations or other pieces of retrieved information Xu et al. (2024b). In high-stakes domains such as healthcare and law, even minor inconsistencies can lead to significant consequences—for instance, conflicting drug dosages from different clinical guidelines (Dayton, 2012; Zhao et al., 2025), or divergent legal interpretations across jurisdictions (Guha et al., 2023).
Recent research proposes techniques such as post-retrieval contradiction detection (Xu et al., 2024a), and confidence-aware re-ranking (Ren et al., 2025) to address these issues. Some frameworks, like MedRAG(Zhao et al., 2025), apply weighted retrieval and ensemble voting to prioritize reliable sources. Others explore neural symbolic consistency checking, where injected knowledge is aligned to pre-defined ontologies or verified using structured reasoning paths (Ciatto et al., 2024).
Another growing area involves alignment-aware reranking, where retrieved documents are filtered based on their alignment with the LLM’s intermediate beliefs (Jin et al., 2025). Future directions may include interactive consistency resolution (e.g., user-in-the-loop conflict selection), as well as integrating factual calibration modules (Dong et al., 2022) that explicitly monitor factuality during decoding. These methods collectively aim to make knowledge-enhanced LLMs more robust, explainable, and reliable in dynamic or sensitive environments.
B.2 Cross-Domain Knowledge Transfer
Cross-domain transfer is a central challenge in building generalized yet specialized LLMs. As LLMs are exposed to knowledge from diverse domains, they must navigate incompatible ontologies, varied domain languages, and distinct reasoning structures Li et al. (2025b). For instance, transferring concepts from chemistry to healthcare involves not only bridging terminology gaps but also adapting to different causal assumptions and data formats (Schroeder et al., 2018).
Several strategies have been proposed to manage this complexity. Adapter-based modularization He et al. (2021) allows domain-specific components to be trained separately and selectively activated. Meta-learning approaches(Hou et al., 2022) help models rapidly adapt to new domains with minimal supervision. Additionally, continual pretraining on mixed-domain corpora (Jin et al., 2022) offers a scalable method to improve robustness without catastrophic forgetting.
Standardized datasets such as CrossNER(Liu et al., 2021) for multilingual named entity recognition, MultiLexSum(Shen et al., 2022) for cross-domain summarization, and MEDIQA-QA (Yadav et al., 2021) for biomedical QA serve as valuable testbeds for cross-domain evaluation. Future work may explore retrieval-augmented transfer, where dynamic selection of domain-relevant knowledge supports adaptive reasoning, or domain-invariant embedding learning, enabling LLMs to generalize across tasks without explicit supervision.
B.3 Scalability and Efficiency of Knowledge Integration
As LLMs are increasingly augmented with large-scale external knowledge such as entire knowledge graphs, document corpora, or real-time retrieval APIs, the computational and memory cost of incorporating this knowledge becomes a bottleneck. Efficient integration remains a key challenge, especially when operating under low-resource or latency-constrained settings.
Techniques such as sparse retrieval (Lee et al., 2019), memory compression (Zhong et al., 2024), and caching strategies have been proposed to reduce overhead. Modular architectures (e.g., adapters or plug-in modules) also allow partial activation of knowledge, improving scalability. Future research could explore task-aware pruning of external knowledge, knowledge distillation from retrieval-based pipelines into compact models, and efficient routing mechanisms to select only relevant knowledge for each input.
B.4 Evaluation and Hallucination Detection
Evaluating knowledge-enhanced LLMs remains difficult due to the lack of standardized benchmarks and automatic metrics for factual consistency, coverage, and reasoning depth. Moreover, LLMs often hallucinate facts even when augmented with accurate knowledge (Ji et al., 2023), making it hard to trust outputs in high-stakes tasks.
Recent work explores metrics like FactScore (Min et al., 2023), entailment-based verification (Patwa et al., 2022), and human-in-the-loop evaluation schemes. However, few of these methods scale across domains or languages. There is a growing need for task-specific, fine-grained evaluation metrics that capture whether the model used the retrieved knowledge effectively and truthfully. Additionally, incorporating hallucination detection as an internal module, such as through consistency checks between generation and source knowledge, may help reduce risk and improve interoperability.